OpenStudy (mikehowk57):

If y is directly proportional to x and y = 3 when x = 18, what is the value of x when y = 20?

OpenStudy (kittiwitti1):

What do you think "directly proportional" means?

OpenStudy (brainzonly):

I cannot answer this myself, but this might help: http://www.mathsisfun.com/algebra/directly-inversely-proportional.html

OpenStudy (kittiwitti1):

^ That is a good link. I would suggest reading that @mikehowk57

OpenStudy (mikehowk57):

isnt that y=kx?

OpenStudy (brainzonly):

Hello, kitti.

OpenStudy (mikehowk57):

i know how to plug it in it just confuses me a bit..

OpenStudy (kittiwitti1):

Hello brainz. How are you ?

OpenStudy (mikehowk57):

would it be.. .16 = k? could you check

OpenStudy (kittiwitti1):

@mikehowk57 Directly proportional means as one goes up, so does the other. Inversely would mean as one goes up the other goes DOWN.

OpenStudy (mikehowk57):

3 = 18k is the thing

OpenStudy (kittiwitti1):

Yup. So what is k?

OpenStudy (mikehowk57):

.16? or .17

OpenStudy (kittiwitti1):

Since \(3\lt18\) it is most likely a decimal. .17 if you round, yes

OpenStudy (mikehowk57):

oh ok thank you

OpenStudy (kittiwitti1):

No problem. :)

OpenStudy (mikehowk57):

then what? do you divide the .17 with the 20? or what?

OpenStudy (kittiwitti1):

You have k=0.17 in the formula y=kx, right?

OpenStudy (kittiwitti1):

Now they give you y=20, so they want the x.

OpenStudy (kittiwitti1):

\[y=kx\]\[y=0.17x\]\[20=0.17x\]\[x=?\]

OpenStudy (brainzonly):

Sorry I didn't answer sooner :P I am fine and taking better care to check definitions LOL How are you?

OpenStudy (kittiwitti1):

Chemistry is kicking my butt, but I am fine too haha

OpenStudy (kittiwitti1):

@mikehowk57 you good?

OpenStudy (brainzonly):

Chemistry is fun! :D Ask me anything but math..

OpenStudy (kittiwitti1):

Lol. I do need help in Chemistry, I suppose that would work!

OpenStudy (kittiwitti1):

Or maybe I keep miscalculating my work. Ew.

OpenStudy (brainzonly):

Not fun D:

OpenStudy (kittiwitti1):

Yeah. :(

OpenStudy (kittiwitti1):

@mikehowk57 How are you doing -- do you need help? Are you stuck?

OpenStudy (brainzonly):

I think he might be trying to solve his own problem.

OpenStudy (mikehowk57):

I got it lol sorry!

OpenStudy (brainzonly):

Good for you!

jhonyy9 (jhonyy9):

y = 3 so x=18 so x= 6*3 y = 20 => x = 6*20 = 120 i think this is the right way to solve this exercise

jhonyy9 (jhonyy9):

hope helped

jhonyy9 (jhonyy9):

kitty your opinion about this above wrote please ?

jhonyy9 (jhonyy9):

ty

OpenStudy (mikehowk57):

The projected profit (in thousands of dollars) of a footwear manufacturing company is modeled by the function f(x) = x2 – 2x – 8 where x is the number of pairs of shoes produced (in thousands). What is the minimum number of shoes that the company needs to produce to avoid a loss?

OpenStudy (kittiwitti1):

@jhonyy9 I think I already said several times, plus it's in my bio, that I do not like being called kitty. ☻

jhonyy9 (jhonyy9):

ohh really sorry but than how ? please << kitti >> will be right ,accepted ?

OpenStudy (kittiwitti1):

yes @jhonyy9 that is OK

jhonyy9 (jhonyy9):

ok. kitti nice to meet you and welcome form Europe

jhonyy9 (jhonyy9):

so where is now in the night 2,10 oclock

jhonyy9 (jhonyy9):

from Europe sorry so this mean that i m tired i will go sleeping now good luck bye bye